OpenStudy (anonymous):

consider the piecewise h(x) = -2/3(x+1) x<-1 -4x-4 -1≤ x≤ 0 x-8 1

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

You ARE going to have to graph those three lines. Just cross out the parts you don't need.

OpenStudy (anonymous):

http://awesomescreenshot.com/0303lr6g9d

OpenStudy (tkhunny):

What do you think? Did we cover anything below y = -6?

OpenStudy (anonymous):

i think [-7, negative infinity)

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

are the pieces continuous at the joints?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

http://awesomescreenshot.com/0303lr6g9d

OpenStudy (amistre64):

well it cant go to negative infinity since there is nothing below -7

OpenStudy (amistre64):

id say -7 and up works fine

OpenStudy (anonymous):

so its [-7, positive infinity)

OpenStudy (amistre64):

-2/3(x+1) x<-1 -4x-4 -1≤ x≤ 0 x-8 1<x<4 -2/3(x+1) -2(x+1)/3 > 0 -4x-4 0>= -4x-4 >= -4 x-8 -7<x-8<-4

OpenStudy (amistre64):

-7 to -4 to 0 to inf

OpenStudy (anonymous):

ah okay so [-7,-4] and [0, positive infinity)

OpenStudy (amistre64):

-4 to 0 is still in the range

OpenStudy (anonymous):

wait so now I'm confused.. how to write it in interval notation? i thought it was [-7, positive infinity)

OpenStudy (amistre64):

it is

OpenStudy (anonymous):

oh okay

OpenStudy (amistre64):

wait, didnt see the lack of equailty on the -7 (-7,inf)

OpenStudy (anonymous):

but it includes 7 sp its [-7, pos inf)

OpenStudy (amistre64):

how does it include -7?

OpenStudy (anonymous):

includes -7

OpenStudy (anonymous):

on the graph of h(x) = x-8

OpenStudy (anonymous):

http://awesomescreenshot.com/0303lr6g9d

OpenStudy (amistre64):

x-8 1<x<4 , x is never equal to 1

OpenStudy (anonymous):

oh my bad. thanks!

OpenStudy (amistre64):

good luck